OpenStudy (anonymous):

WILL MEDAL AND FAN!! Simplify. 2 + (–10). –12 12 8 –8

OpenStudy (itsbribro):

-8

OpenStudy (anonymous):

-8 because if you flip the problem around, -10+2= -8

OpenStudy (itsbribro):

exactly what @Htester said i was going to say that

OpenStudy (anonymous):

\[2+(-10)=2-10=-8\]

OpenStudy (itsbribro):

just think of it this way if you have -10 $ in the bank account and your adding 2 dollars to the account it would be -8